# DESI Data Release 1: Stellar Catalogue

### v1

@@ Abstract
In this paper we present the stellar Value-Added Catalogue (VAC) based on the DESI Data Release 1. This VAC contains stellar parameter, abundance and radial velocity measurements for more than 4 million stars. It also contains, for the first time, measurements from individual epochs for more than a million stars with at least two observations. The main contribution to the catalogue comes from the bright program of the main survey, which includes 2.5 million stars, and the backup program, which includes 1 million stars. The combined magnitude range for the stars in the catalogue extends from Gaia G to G . For the magnitude range this catalogue represents a factor of 10 increase in the number of stars with radial velocity and abundance measurements compared to existing surveys. Despite DESI’s resolution (R ), the median radial velocity random error for stars in the catalogue is better than 1 km s-1. The stellar parameters and abundances of stars in DESI are measured by two independent pipelines, and after applying a temperature-dependent calibration, [Fe/H] abundances of high signal-to-noise stars are accurate to better than 0.1 dex when compared to high-resolution surveys. The catalogue probes different Galactic components including a particularly large number of distant stars: tens of thousands of stars further than 10 kpc, and thousands further than 50 kpc. The catalogue also contains several thousand extremely metal-poor stars with . The released sample of stars includes measurements for thousands of stars that are members of dwarf galaxies, open and globular clusters as well as members of several dozen stellar streams. The next public DESI data release is expected in two years and will contain three times as many stars as DR1.
@@ 1 Introduction
The first data set from DESI was released publicly as the Early Data Release (EDR) (DESI Collaboration et al., 2024c). Those data were based on survey validation (DESI Collaboration et al., 2024b) observations and amounted to some 1.5 million spectra, including about 400,000 stellar spectra. The analysis of the stellar subset was presented in Koposov et al. (2024b), K24 hereafter, showing that DESI is capable of measuring radial velocities accurate to km s-1 and chemical abundances to dex. These data have already been used in several studies Manser et al. (2024); Valluri et al. (2025).
The DESI collaboration recently made publicly available the Data Release 1 (DR1) (DESI Collaboration et al., 2025) containing more than 18 million redshifts and approximately four million stellar spectra observed during the first 13 months of main survey operations. These data have been used for multiple cosmological analyses, such as BAO and clustering measurements using different tracers leading to new constraints on cosmological parameters (DESI Collaboration et al., 2024a). This paper presents the analysis of the stellar subset of the DESI DR1 and an associated Value-Added Catalogue (VAC). In the paper, we highlight the improvements over previous releases and provide guidance on the use of this catalogue. The paper is structured as follows. In Section 2 we describe the key processing steps and improvements that went into the Value-Added Catalogue. In Section 3 we summarise the contents of the Catalogue. In Section 4 we validate the abundance and velocity measurements. In Section 5 we discuss the possible scientific uses of this dataset as well as compare it to other existing large spectroscopic surveys.
@@ 2 Data
The DESI instrument is a fiber-fed spectrograph that is capable of observing 5,000 spectra simultaneously over an 8-square-degree field-of-view. The DESI wavelength coverage extends from 3600 Å in the blue to 9800 Å in the red. The spectra of each object are split by the dichroic beam splitter into three spectrographs that observe the wavelength ranges from 3600 Å to 5930 Å (B arm), 5600 Å to 8000 Å (R arm), and 7470 Å to 9800 Å (Z arm). The average resolution of the spectra varies with wavelength from at 3600 AA to R at 9800 Å. All the DESI spectra are extracted at fixed sampling of 0.8 Å per pixel in the barycentric frame using spectral perfectionism framework (Bolton & Schlegel, 2010), and are flux-calibrated using spectra of stars observed in each exposure.
DESI data are organised by surveys and programs. Table 1 gives the list of surveys and programs included in the DR1 together with the number of spectra analysed for this Value-Added Catalogue. The surveys sv1, sv2, sv3 correspond to the survey validation programs (DESI Collaboration et al., 2024b), and survey main is the main DESI survey. The survey special corresponds to various dedicated pointings and observations, such as observations of M31 used in Dey et al. (2023). The surveys themselves consist of several programs, such as dark, bright, and backup. The programs are completely independent of each other, have different target selection algorithms, and are triggered under different observing conditions. The dark program is executed in the best conditions. The bright program is executed in less-optimal conditions, such as when the moon is up, while the backup program is conducted under conditions that are below those acceptable for the bright program. The formal separation of conditions into backup, bright and dark is done using the so-called survey speed, which is determined from the seeing, sky background, transparency, and airmass at the time of observations (Schlafly et al., 2023). The dark program includes the main cosmological DESI survey. The bright program includes the Bright Galaxy Survey (BGS) (Hahn et al., 2023) and the Milky Way Survey (MWS) (Cooper et al., 2023). The backup program is a separate survey itself, described in detail in Dey et al. (2025), and is devoted to observations of bright stars with an extended footprint compared to the MWS survey.
@@ 2.2 MWS processing
The MWS pipelines do not analyse every spectrum observed by DESI and processed by the DESI pipeline. They instead focus on spectra that are of high enough quality and that could belong to stars irrespective of the survey and program they have been observed in. Specifically for this data release, only spectra satisfying all of the following requirements have been processed111Here we rely on various DESI processing columns defined in Guy et al. (2023) and https://desidatamodel.readthedocs.io/en/latest/ :
@@ 2.3.1 RVS pipeline
For this release we use an updated version of the RVSpecFit code, where the stellar spectrum model is constructed using a neural network emulator. This was done to address multiple problems highlighted in K24. We use a feed-forward fully-connected neural network with four input parameters , , , . There are three 128-d hidden layers of the network, and the SiLU (Sigmoid Linear Unit) is used as a non-linearity. These layers of the network return a 200-d vector which is then passed through a final fully linear layer transforming 200-d vector into a DESI-like spectrum (i.e., spectrum with several thousands pixels, 0.4 Å sampling and spanning the wavelength range of DESI data). The neural network was implemented using the pytorch framework and trained on PHOENIX spectra using the Adam optimiser (Kingma & Ba, 2017) with the ReduceOnPlateau scheduler.
@@
Another modification to RVS pipeline is an additional warning flag that is set in the RVS_WARN bitmask column. The fourth bit (corresponding to value of 8) of this bitmask is set if some of the stellar parameters are close or beyond the stellar grid edges. Specifically, if the [Fe/H] is within 0.01 dex of or , and if the is within 10 K of 2300 K or 15,000 K the fourth bit of RVS_WARN is enabled. We recommend that when analysing stellar parameters from RVS, only stars with RVS_WARN=0 should be used, but at the same time many stars that have well measured radial velocities and are well fit by RVS have this warning flag.
The change of the interpolator is the main improvement to the RVS pipeline with respect to the version used for the construction of the EDR VAC. Other (small) improvements and bug fixes were also adopted, such as addressing the issues with stars with high rotational velocity, , described in the EDR paper, and improving the uncertainties of stellar parameters.
@@ 2.3.2 SP pipeline
The SP pipeline is powered by FERRE (Allende-Prieto & Apogee Team, 2023) code that is designed to measure stellar atmospheric parameters and individual elemental abundances by fitting continuum normalised spectra. There were no substantial changes in the SP pipeline, other than an increase in the number of elements for which abundances are derived, from four to ten. It now includes: Fe, Ca, C, Mg, Si, Na, Cr, Ni, Al, and Ti, reported in that order in the ELEM array within the SPTAB extension of the summary files. As with the EDR data, the results that were fitted with Kurucz model atmospheres (Kurucz, 2005) should be used, avoiding those fit with PHOENIX models. This means we suggest to use BESTGRID ‘s_rdesi1’ when working with SP measurements. Despite the efforts to include more chemical elements, we deem only the elemental abundances of Mg, Ca, and Fe to be reliable.
@@ 3 Contents of the release
The file structure of the catalogue and column names have been described in K24, and are mostly identical in this VAC. The data model is available online222https://desi-mws-dr1-datamodel.readthedocs.io/en/latest/.
Similar to the EDR VAC, this release includes measurements by RVS and SP pipelines of coadded spectra from different surveys and programs. The release also contains a merged catalogue (named mwsall-pix-iron.fits) that contains the combination of the RVS and SP measurements across all the surveys, targeting information, and the crossmatch with Gaia DR3 source catalogue (Gaia Collaboration et al., 2021, 2023). This combined catalogue is probably the most useful to work with for researchers interested in Milky Way science with DESI. We provide a short reminder of the structure of this file in the Appendix A, and also release a tutorial on how to use the catalogue 333https://github.com/desimilkyway/dr1_tutorials/.
@@ 3.1 Co-added spectra analysis
The numbers in Table 1 may include duplicates, as objects observed in multiple programs are counted for each. The combined catalogue of measurements (the mwsall file) for the stellar VAC includes measurements from all the surveys and programs, and in the case of objects observed in more than one program and survey, the observation with the highest signal-to-noise (in the R arm) is marked as primary (this can be identified using the PRIMARY column in the table)444The identification of primary observation happens not only for sources that have the same TARGETID in multiple programs and surveys, but also for sources with different TARGETID that are found to be within 0.5 arc-seconds of each other at epoch of J2021.0. To calculate the source position at J2021 we use the proper motions from the targeting catalogue FIBERMAP.. The number of primary (unique) objects in the combined catalogue is 5,930,124. 4,378,110 of those are classified as a star by Redrock
@@ 3.1.1 Survey footprint
The backup program footprint extends further than the other programs and beyond the DR9 footprint of the DECaLS survey (Dey et al., 2019), as the backup targets are selected using Gaia rather than DECaLS photometry (see Dey et al., 2025, for more details). The figure shows that the density of targets in the catalogue varies dramatically across the sky for dark and bright programs, because of the different degree of progress of the survey in different areas of the sky. The reason for this is that the bright program requires five passes across the survey footprint, while the for the dark program seven passes are required (see Schlafly et al., 2023). The backup program is done using a single pass over the footprint, which results in a more uniform density of objects in the catalogue seen in Figure 1 (roughly equal to DESI’s fiber density of 5000 objects per 8 square degree of field of view). The dark program has a significantly lower density of stellar targets compared to the bright or backup programs, as the majority of observed stars in the dark program are only observed as flux calibration standards (which typically occupy some 100 fibres out of 5000 – see Guy et al. 2023).
@@ 3.1.2 Survey targets
As described earlier, the DESI stellar catalogue contains objects selected in various ways in different programs. However the dominant source of stars in it is the Milky Way Survey. The MWS targeting strategy has been described in Myers et al. (2023) and Cooper et al. (2023), but here we highlight the actual targets observed and included in the VAC, while focusing on targets observed in the main survey.Figure 2 shows the colour-magnitude diagrams for objects classified as stars in three programs of the main survey and the bright program of the survey validation 3 (sv3) for comparison. The magnitudes were not corrected for extinction in this plot. The colour-magnitude distribution of stars in the bright program of the main survey is mostly showing the target selections of the MWS and is similar to the colour-magnitude distribution in the sv3 survey, as little was changed in the selection. The distribution shows that the stellar targets in the bright survey are mostly limited to , with a small fraction of fainter sources. The colour-magnitude diagram for the dark program of the main survey shows that most objects have magnitudes similar to the bright program, but they are within a narrow colour range around 0.3. This reflects the selection of standards for spectro-photometric calibration of DESI. The objects at fainter magnitudes () in the dark program are either high priority MWS targets observed in both the dark and bright surveys, or additional secondary targets (such as faint blue horizontal branch (BHBs), see Byström et al., 2024).
@@
While the discussion of the survey selection function is beyond the scope of this paper, we stress that the input targets catalogues, as well their selection criteria for the main survey were published (Myers et al., 2023) and are available at https://data.desi.lbl.gov/public/dr1/target/. Within a given survey and program (e.g. bright program of the main survey), we expect that for objects that are only members of one target class (i.e. MWS_MAIN_BLUE) the probability of having a DESI spectrum is only a function the number of DESI passes over given area and possibly local target density, but not any of the source properties such as magnitude, colour or proper motion.
@@ 3.2 Single-epoch measurements
Table 3 gives a total number of Gaia sources in this VAC with more than one DESI observation within different surveys and programs (note that some sources have been observed many times across multiple surveys and programs as well). The table shows that the largest number of repeated observations comes from the backup program. This is caused by the conditions of observations being poor and DESI tiles in that program often requiring more than one exposure. Around half a million objects have also been observed multiple times in the bright program of the main survey, followed by quarter of a million of objects with re-observations in the dark program. The sv3 survey also provides a significant number of objects with repeated measurements, because of the so-called 1% survey (see DESI Collaboration et al., 2024b), where a small area of the sky was observed repeatedly to achieve high degree of completeness across all target classes.
The total number of single-epoch spectra measurements for this VAC is 10,012,925. This includes around 4 million spectra for objects observed only once, where the spectrum should be identical to the coadded spectrum.
In Figure 4 we show the distribution of the number of repeated DESI observations for different Gaia objects, grouped by the maximum interval between earliest and latest observations. The figure shows that most objects that were observed more than once have a small number (2 to 3) of repeated observations, but there is an extended tail to many tens of observations, with more than objects having ten or more repeats. Most of the objects with a large number of repeated observations have been observed in the survey validation program. The figure also shows that while most objects with repeated observations were observed within a short time span, a substantial fraction of objects have observations separated by a few months. Note that this VAC is based on the dataset obtained only over 13 months of the main survey and 5 months of survey validation and since then DESI has been observing for three years, the future data release should have a significantly larger baseline of repeated observations.

### v3

@@ Abstract
In this paper we present the stellar Value-Added Catalogue (VAC) based on the DESI Data Release 1. This VAC contains stellar parameter, abundance and radial velocity measurements for more than 4 million stars. It also contains, for the first time, measurements from individual epochs for more than a million stars with at least two observations. The main contribution to the catalogue comes from the bright program of the main survey, which includes 2.5 million stars, and the backup program, which includes 1 million stars. The combined magnitude range for the stars in the catalogue extends from Gaia G to G . For the magnitude range this catalogue represents a factor of 10 increase in the number of stars with radial velocity and abundance measurements compared to existing surveys. Despite DESI’s resolution (R ), the median radial velocity uncertainty for stars in the catalogue is better than 1 km s-1. The stellar parameters and abundances of stars in DESI are measured by two independent pipelines, and after applying a temperature-dependent calibration, [Fe/H] abundances of high signal-to-noise stars are accurate to better than 0.1 dex when compared to high-resolution surveys. The catalogue probes different Galactic components including a particularly large number of distant stars: tens of thousands of stars further than 10 kpc, and thousands further than 50 kpc. The catalogue also contains several thousand extremely metal-poor stars with . The released sample of stars includes measurements for thousands of stars that are members of dwarf galaxies, open and globular clusters as well as members of several dozen stellar streams. The next public DESI data release is expected in less than two years and will contain three times as many stars as DR1.
@@ 1 Introduction
The first data set from DESI was released publicly as the Early Data Release (EDR) (DESI Collaboration et al., 2024b). Those data were based on survey validation (DESI Collaboration et al., 2024a) observations and amounted to some 1.5 million spectra, including about 400,000 stellar spectra. The analysis of the stellar subset was presented in Koposov et al. (2024a), K24 hereafter, showing that DESI is capable of measuring radial velocities accurate to km s-1 and chemical abundances to dex. These data have already been used in several studies (Manser et al., 2024; Valluri et al., 2025).
The DESI collaboration recently made publicly available the Data Release 1 (DR1) (DESI Collaboration et al., 2025) containing more than 18 million redshifts and approximately four million stellar spectra observed during the first 13 months of main survey operations. These data have been used for multiple cosmological analyses, such as BAO and clustering measurements using different tracers leading to new constraints on cosmological parameters (Adame et al., 2025). This paper presents the analysis of the stellar subset of the DESI DR1 and an associated Value-Added Catalogue (VAC). In the paper, we highlight the improvements over previous releases and provide guidance on the use of this catalogue. The paper is structured as follows. In Section 2 we describe the key processing steps and improvements that went into the Value-Added Catalogue. In Section 3 we summarise the contents of the Catalogue. In Section 4 we validate the abundance and velocity measurements. In Section 5 we discuss the possible scientific uses of this dataset as well as compare it to other existing large spectroscopic surveys.
@@ 2 Data
The DESI instrument is a fiber-fed spectrograph that is capable of observing 5,000 spectra simultaneously over an 8-square-degree field-of-view. The DESI wavelength coverage extends from 3600 Å in the blue to 9800 Å in the red. The spectra of each object are split by the dichroic beam splitter into three spectrographs that observe the wavelength ranges from 3600 Å to 5930 Å (B arm), 5600 Å to 8000 Å (R arm), and 7470 Å to 9800 Å (Z arm). The average resolution of the spectra varies with wavelength from at 3600 Å to R at 9800 Å. All the DESI spectra are extracted at fixed sampling of 0.8 Å per pixel in the barycentric frame using spectral perfectionism framework (Bolton and Schlegel, 2010), and are flux-calibrated using spectra of stars observed in each exposure.
DESI data are organised by surveys and programs. Table 1 gives the list of surveys and programs included in the DR1 together with the number of spectra analysed for this Value-Added Catalogue. The surveys sv1, sv2, sv3 correspond to the survey validation programs (DESI Collaboration et al., 2024a), and survey main is the main DESI survey. The survey special corresponds to various dedicated pointings and observations, such as observations of M31 used in Dey et al. (2023). The surveys themselves consist of several programs, such as dark, bright, and backup. The programs are completely independent of each other, have different target selection algorithms, and are triggered under different observing conditions. The dark program is executed in the best conditions. The bright program is executed in less-optimal conditions, such as when the moon is up, while the backup program is conducted under conditions that are below those acceptable for the bright program. The formal separation of conditions into backup, bright and dark is done using the so-called survey speed, which is determined from the seeing, sky background, transparency, and airmass at the time of observations (Schlafly et al., 2023). The dark program includes the main cosmological DESI survey. The bright program includes the Bright Galaxy Survey (BGS) (Hahn et al., 2023) and the Milky Way Survey (MWS) (Cooper et al., 2023). The backup program is a separate survey itself, described in detail in Dey et al. (2025), and is devoted to observations of bright stars with an extended footprint compared to the MWS survey.
@@ 2.2 MWS processing
The MWS pipelines do not analyse every spectrum observed by DESI and processed by the DESI pipeline. They instead focus on spectra that are of high enough quality and that could belong to stars irrespective of the survey and program they have been observed in. Specifically for this data release, only spectra satisfying all of the following requirements have been processed:111Here we rely on various DESI processing columns defined in Guy et al. (2023) and https://desidatamodel.readthedocs.io/en/latest/
@@ 2.3.1 RVS pipeline
For this release we use an updated version of the RVSpecFit code, where the stellar spectrum model is constructed using a neural network emulator. This was done to address multiple problems highlighted in K24. We use a feed-forward fully-connected neural network with four input parameters , , , . There are three 128-d hidden layers of the network, and the SiLU (Sigmoid Linear Unit) is used as a non-linearity. These layers of the network return a 200-d vector which is then passed through a final fully linear layer transforming 200-d vector into a DESI-like spectrum (i.e., spectrum with several thousands pixels, 0.4 Å sampling and spanning the wavelength range of DESI data). The neural network was implemented using the pytorch framework and trained on PHOENIX spectra using the Adam optimiser (Kingma and Ba, 2017) with the ReduceLROnPlateau scheduler.
@@
Another modification to RVS pipeline is an additional warning flag that is set in the RVS_WARN bitmask column. The fourth bit (corresponding to value of 8) of this bitmask is set if some of the stellar parameters are close to or beyond the stellar grid edges. Specifically, if the [Fe/H] is within 0.01 dex of or , and if the is within 10 K of 2,300 K or 15,000 K the fourth bit of RVS_WARN is enabled. We recommend that when analysing stellar parameters from RVS, only stars with RVS_WARN=0 should be used, but at the same time many stars that have well measured radial velocities and are well fit by RVS have this warning flag.
The change of the interpolator is the main improvement to the RVS pipeline with respect to the version used for the construction of the EDR VAC. Other small improvements and bug fixes were also adopted, such as addressing the issues with stars with high rotational velocity, , described in the EDR paper, and improving the uncertainties of stellar parameters.
@@ 2.3.2 SP pipeline
The SP pipeline is powered by FERRE (Allende-Prieto and Apogee Team, 2023) code that is designed to measure stellar atmospheric parameters and individual elemental abundances by fitting continuum normalised spectra. There were no substantial changes in the SP pipeline, other than an increase in the number of elements for which abundances are derived, from four to ten. It now includes: Fe, Ca, C, Mg, Si, Na, Cr, Ni, Al, and Ti, reported in that order in the ELEM array within the SPTAB extension of the summary files. As with the EDR data, the results that were fitted with Kurucz model atmospheres (Kurucz, 2005) should be used, avoiding those fitted with PHOENIX models. This means we suggest to use BESTGRID ‘s_rdesi1’ when working with SP measurements. Despite the efforts to include more chemical elements, we deem only the elemental abundances of Mg, Ca, and Fe to be reliable.
@@ 3 Contents of the release
The file structure of the catalogue and column names have been described in K24, and are mostly identical in this VAC. The data model is available online.222https://desi-mws-dr1-datamodel.readthedocs.io/en/latest/
Similar to the EDR VAC, this release includes measurements by RVS and SP pipelines of coadded spectra from different surveys and programs. The release also contains a merged catalogue (named mwsall-pix-iron.fits) that contains the combination of the RVS and SP measurements across all the surveys, targeting information, and the crossmatch with Gaia DR3 source catalogue (Gaia Collaboration et al., 2016, 2021, 2023). This combined catalogue is probably the most useful to work with for researchers interested in Milky Way science with DESI. We provide a short reminder of the structure of this file in the Appendix A, and also release a tutorial on how to use the catalogue.333https://github.com/desimilkyway/dr1_tutorials/ The validation of the data quality of the catalogue is discussed in Section 4, but the full list of known issues affecting the catalogue that users need to be aware of is provided in the Appendix B.
@@ 3.1 Co-added spectra analysis
The numbers in Table 1 may include duplicates, as objects observed in multiple programs are counted for each. The combined catalogue of measurements (the mwsall file) for the stellar VAC includes measurements from all the surveys and programs, and in the case of objects observed in more than one program and survey, the observation with the highest signal-to-noise (in the R arm) is marked as primary (this can be identified using the PRIMARY column in the table).444The identification of primary observation happens not only for sources that have the same TARGETID in multiple programs and surveys, but also for sources with different TARGETID that are found to be within 0.5 arc-seconds of each other at epoch of J2021.0. To calculate the source position at J2021 we use the proper motions from the targeting catalogue FIBERMAP. The number of primary (unique) objects in the combined catalogue is 5,930,124. 4,378,110 of those are classified as a star by Redrock
@@ 3.1.1 Survey footprint
The backup program footprint extends further than the other programs and beyond the DR9 footprint of the DECaLS survey (Dey et al., 2019), as the backup targets are selected using Gaia rather than DECaLS photometry (see Dey et al., 2025, for more details). The figure shows that the density of targets in the catalogue varies dramatically across the sky for dark and bright programs, because of the different degree of progress of the survey in different areas of the sky. The reason for this is that the bright program requires five passes across the survey footprint, while for the dark program seven passes are required (see Schlafly et al., 2023). The backup program is done using a single pass over the footprint, which results in a more uniform density of objects in the catalogue seen in Figure 1 (roughly equal to DESI’s fiber density of 5000 objects per 8 square degree of field of view). The dark program has a significantly lower density of stellar targets compared to the bright or backup programs, as the majority of observed stars in the dark program are only observed as flux calibration standards (which typically occupy some 100 fibres out of 5000 – see Guy et al. 2023).
@@ 3.1.2 Survey targets
As described earlier, the DESI stellar catalogue contains objects selected in various ways in different programs. However the dominant source of stars in it is the Milky Way Survey. The MWS targeting strategy has been described in Myers et al. (2023) and Cooper et al. (2023), but here we highlight the actual targets observed and included in the VAC, while focusing on targets observed in the main survey. Figure 2 shows the colour-magnitude diagrams for objects classified as stars in three programs of the main survey and the bright program of the survey validation 3 (sv3) for comparison. The magnitudes were not corrected for extinction in this plot. The colour-magnitude distribution of stars in the bright program of the main survey is mostly showing the target selections of the MWS and is similar to the colour-magnitude distribution in the sv3 survey, as little was changed in the selection. The distribution shows that the stellar targets in the bright survey are mostly limited to , with a small fraction of fainter sources. The colour-magnitude diagram for the dark program of the main survey shows that most objects have magnitudes similar to the bright program, but they are within a narrow colour range around 0.3. This reflects the selection of standards for spectro-photometric calibration of DESI. The objects at fainter magnitudes () in the dark program are either high priority MWS targets observed in both the dark and bright surveys, or additional secondary targets (such as faint blue horizontal branch (BHBs), see Byström et al., 2025).
@@
While the discussion of the survey selection function is beyond the scope of this paper, we stress that the input target catalogues, as well as their selection criteria for the main survey were published (Myers et al., 2023) and are available at https://data.desi.lbl.gov/public/dr1/target/. Within a given survey and program (e.g. bright program of the main survey), we expect that for objects that are only members of one target class (i.e. MWS_MAIN_BLUE) the probability of having a DESI spectrum is only a function of the number of DESI passes over given area and possibly local target density, but not any of the source properties such as magnitude, colour or proper motion.
@@ 3.2 Single-epoch measurements
Table 3 gives a total number of Gaia sources in this VAC with more than one DESI observation within different surveys and programs (note that some sources have been observed many times across multiple surveys and programs as well). The table shows that the largest number of repeated observations comes from the backup program. This is caused by the conditions of observations being poor and DESI tiles in that program often requiring more than one exposure. Around half a million objects have also been observed multiple times in the bright program of the main survey, followed by quarter of a million objects with re-observations in the dark program. The sv3 survey also provides a significant number of objects with repeated measurements, because of the so-called 1% survey (see DESI Collaboration et al., 2024a), where a small area of the sky was observed repeatedly to achieve high degree of completeness across all target classes.
The total number of single-epoch spectra measurements for this VAC is 10,012,925. This includes around 4 million spectra for objects observed only once, where the spectrum is mostly identical to the coadded spectrum.555The coadded single epoch spectra are not exactly identical to single epoch spectra due to an identified resolution matrix and masking bug that affects a tiny fraction of pixels in DESI spectra, see Appendix E in DESI Collaboration et al. (2025)
In Figure 4 we show the distribution of the number of repeated DESI observations for different Gaia objects, grouped by the maximum interval between earliest and latest observations. The figure shows that most objects that were observed more than once have a small number (2 to 3) of repeated observations, but there is an extended tail to many tens of observations, with more than objects having ten or more repeats. Most of the objects with a large number of repeated observations have been observed in the survey validation program. The figure also shows that while most objects with repeated observations were observed within a short time span, a substantial fraction of objects have observations separated by a few months. Note that this VAC is based on the dataset obtained only over 13 months of the main survey and 5 months of survey validation. Since then DESI has been observing for three years, so the future data release will have a significantly larger baseline of repeated observations.